Richards age in 11 years will be three times what it was 15 years ago. How old is Richard now?
Answer by fmo(8)

Let's let x represent Richard's age.
Richard's age in 11 years, can then be represented as x + 11.
15 years ago, Richard's age would have been x - 15.
Finally, we know that Richard's age in 11 years (x + 11) will be three times what it was 15 years ago (3 * (x - 15)), and that forms an equation that looks like:
x + 11 = 3 * (x - 15)
Now to solve...
x + 11 = 3(x - 15) (problem restated)
x + 11 = 3x - 45 (right side expanded)
11 = 2x - 45 (subtract x from both sides)
56 = 2x (add 45 to both sides)
28 = x (divide both sides by 2)
Richard is 28 years old.